# The Lunar Gravitational-wave Antenna: Mission Studies and Science Case

### v1

@@ 3.3.2 Supernovae
Main contributors: Ferdinando Patat, Stefano Benetti, Marica Branchesi, Enrico Cappellaro, Elisabeth-Adelheid Keppler, Francesco Longo, Sourav Roy Chowdhury, David Vartanyan
@@
Thermonuclear explosions, better known as Type Ia Supernovae (SN Ia), play a crucial role in modern cosmology. In addition to providing the first evidence that the expansion of the Universe is accelerated by some sort of dark energy, they recently led to a tension with the value of the Hubble constant derived from the Cosmic Microwave Background (CMB) modelling (see [821] for a review). Despite their importance, the nature of their progenitors is still unclear and two alternative scenarios have been proposed: the SD scenario and the DD scenario (see Sec. 3.2.2). The pieces of evidence collected so far in favour of one or the other option are all indirect. As examples of relevant results, we can list: i) the presence [822] or absence [823] of signatures of circumstellar gas surrounding the progenitor that favour SD and DD, respectively; ii) the failed search for a surviving companion of historical SN Ia in our Galaxy, favoring DD [824]; iii) the failed search of a precursor star in pre-explosion archival images [825] favoring DD; iv) the search for candidate DD systems that did not find a sufficient number of short period, massive WDs to account for the observe SN Ia rates [781] favoring SD, or otherwise sub-Chandrasekhar explosions; v) the distribution of delay time from star formation to explosion derived from statistics of SN Ia rate that tends to favor the DD scenario [826], although the strength of this argument is debatable [827].
@@
Arguably, the ultimate proof of the DD-SNIa connection would be the observation of the merging event followed by the detection of its electromagnetic (EM) counterpart. However, while in some theoretical scenarios the SN explosion and hence the appearance of the EM counterpart, follows a few seconds after the merger [814], others allow for a long delay between the merger and the explosion, extending to $10^{4}$ yr [830]. This implies that, even when a DWD merger with the appropriate mass occurs and its GW emission is detected, it may not be associated with a prompt SN Ia EM event. Hence, it would not provide the long-sought and much-desired smoking gun. This is why, in the current state of affairs, the comparison between the known rate of SNe Ia and the rate of GW sources with the properties expected for merging DD systems is probably a more reliable tool to answer this key question (as described in Sect 3.2.2).
Interestingly enough, the simultaneous detection of a “peculiar” SN Ia in both GW and EM domains is expected for thermonuclear events arising from WD-WD systems made by a CO-WD and a He-WD. These objects should explode via the detonation of a thin He-layer piled-up on the CO-WD via stable mass transfer driven by GW emission with maximum expected frequency of several mHz (see [831] for details). Multi-messenger astronomy could help determining the contribution of double-detonating WDs to the observed SNe Ia rate and, more broadly, allow one to infer the chirp masses for some of the transient, hence constraining the nature of the progenitor. As shown in Sec. 2.2.1, LGWA can observe DWD signals out to several tens of Mpc, which means that a few multi-messenger observations of these sources per year are conceivable.
@@
Following the gravitational collapse of a massive star ($M\geq 8M_{\odot}$), the inner core rebounds and collides with the infalling stellar material producing a shock wave (see [832] for a general review). As the simulations show, the shock wave tends to stall within the infalling material. Extensive theoretical work done in recent years has demonstrated that it is the neutrino heating coming from the accreting proto-NS (coupled to the neutrino-driven turbulent convection between the shock cavity and the stellar surface) which re-powers the shock and leads to subsequent explosion ([833, 324, 315] and references therein). The shock break-out, i.e., the moment when the blast wave reaches the stellar surface, occurs minutes to days post collapse, and shortly after starts emitting electromagnetic radiation, appearing as a SN. The star is destroyed leaving behind a compact remnant.
@@
The LGWA and other future GW detectors will open the possibility of observing these events through all three messengers: EM, GW, neutrinos. Clearly, the simultaneous detection of all three would be the next watershed detection in the field of GWs. Ongoing efforts to understand GW emission from CCSNe over a broad range of frequencies are described in Sec. 3.2.1.
@@ 3.3.3 Tidal Disruption Events
Indeed, ZTF EM follow-up searches of IceCube alerts have been successful in the identification of the TDE AT2019dsg [892] as the optical counterpart of a very high energy neutrino ($\sim$0.2 PeV, IceCube event IC191001A), which represents the first multi-messenger observation of a TDE. Following this remarkably discovery, other two TDE candidates have been associated with two high energy neutrinos events: AT2019fdr (IceCube event IC200530A [893]) and AT2019aalc (IceCube event IC191119A [894]). Interestingly, all the three neutrino-emitting TDEs discovered so far show some common observational features, such as the delayed neutrino signal ($\sim$100 days after the BB peak), a delayed strong dust echoes in the IR and the X-ray detection. This can be ascribed to the physical properties of the system after the star disruption and raise important question on the mechanisms behind the neutrino production itself [895].
@@ 3.3.4 Intermediate and Massive Black Hole Binaries
Where to look for electromagnetic candidates for LGWA emitters ? – The tight correlations observed between the host-galaxy properties and the BH mass indicate that associated EM signals produced from the more massive BHs detected by LGWA, i.e., upwards of 105 – 106 M, should be mainly searched for in dwarf galaxies. Even more challenging is the case of lower-mass BHs (IMBHs). Indeed, besides some indications from GWs [929, 353], any conclusive electromagnetic evidence for the existence of even a single IMBH is still missing. Among the best candidates for accreting IMBHs, we can include ultra-luminous X-ray sources (ULXs; e.g., [930, 931]). ULXs are defined as point-like, off-nuclear (i.e., sources showing offsets from the dynamical center of their hosts) X-ray sources in nearby galaxies; they are characterized by X-ray luminosity above the Eddington limit for a stellar BH (LX$>$1039 erg s-1). In particular, the still-limited (in number) high-mass end of ULX distribution could be powered by IMBHs. ULXs are largely associated with star-forming regions in spiral and irregular galaxies and possibly blue compact dwarf galaxies; however, cases where they trace the old stellar population in elliptical galaxies are also present (see [932]). Since low-mass BHs in shallow potentials (as, e.g., those of dwarf galaxies) are likely to never settle at the center of their galaxies [933], we may expect that the location of IMBHs is consistent with the offset position of ULXs. Another astrophysical environment in which low-mass BHs may form is that of globular clusters [904]. As dense stellar clusters evolve, the compact remnants sink to the center of the cluster and segregate in mass. Wandering IMBHs in the halo of galaxies are also expected as a result of BH-BH encounters in globular clusters that can eject IMBHs. These objects can also represent one of the consequences of the galaxy merging process if seed BHs were common in the first galaxies that merged to form today’s massive galaxies.
@@ 3.3.5 Extreme/Intermediate Mass-ratio Inspirals
So far the signal observed from QPEs is restricted to the soft X-ray band ($E<2$ keV) and the abundance rates of these transients are low ($\approx 10^{-6}\,$pc-3; Arcodia et al., subm.). Therefore, for the purpose of QPE discovery, a soft X-ray telescope with a large effective area and a large field of view is desirable. The WFI onboard Athena may discover $\approx 1$ QPE source per square-degree survey (Arcodia et al., subm.). WFI will also be able to study spectral and timing properties of QPEs deeper in redshift, up to $z\sim 0.5$ for the most common QPE sources peaking at $\sim 10^{42}$ erg s-1 and up to $z\sim 1.3$ for the much rarer ones peaking at $\sim 10^{43}$ erg s-1. The large effective area and high spectral resolution of the X-IFU onboard Athena will also allow to assess the presence of ionized outflowing/inflowing gas before, during, and after QPEs.

### v2

@@ 3.3.2 Supernovae
Main contributors: Ferdinando Patat, Stefano Benetti, Marica Branchesi, Enrico Cappellaro, Kiranjyot Gill, Elisabeth-Adelheid Keppler, Francesco Longo, Sourav Roy Chowdhury, David Vartanyan
@@
Thermonuclear explosions, better known as Type Ia Supernovae (SN Ia), play a crucial role in modern cosmology. In addition to providing the first evidence that the expansion of the Universe is accelerated by some sort of dark energy, they recently led to a tension with the value of the Hubble constant derived from the Cosmic Microwave Background (CMB) modelling (see [838] for a review). Despite their importance, the nature of their progenitors is still unclear and two alternative scenarios have been proposed: the single-degenerate (SD) scenario and the double-degenerate (DD) scenario (see Sec. 3.2.2). The pieces of evidence collected so far in favour of one or the other option are all indirect. As examples of relevant results, we can list: i) the presence [839] or absence [840] of signatures of circumstellar gas surrounding the progenitor that favour SD and DD, respectively; ii) the failed search for a surviving companion of historical SN Ia in our Galaxy, favoring DD [841]; iii) the failed search of a precursor star in pre-explosion archival images [842] favoring DD; iv) the search for candidate DD systems that did not find a sufficient number of short period, massive WDs to account for the observe SN Ia rates [472] favoring SD, or otherwise sub-Chandrasekhar explosions; v) the distribution of delay time from star formation to explosion derived from statistics of SN Ia rate that tends to favor the DD scenario [494], although the strength of this argument is debatable [843].
@@
Arguably, the ultimate proof of the DD-SNIa connection would be the observation of the merging event followed by the detection of its electromagnetic (EM) counterpart. However, while in some theoretical scenarios the SN explosion and hence the appearance of the EM counterpart, follows a few seconds after the merger [505], others allow for a long delay between the merger and the explosion, extending to $10^{4}$ yr [846]. This implies that, even when a DWD merger with the appropriate mass occurs and its GW emission is detected, it may not be associated with a prompt SN Ia EM event. Hence, missing the long-sought and much-desired smoking gun signature. The comparison between the known rate of SNe Ia and the rate of GW sources exhibiting properties expected for merging DD systems does not require the DD-SN Ia association to answer this key question (as described in Sec. 3.2.2). As shown in Sec. 2.2.1, LGWA can observe DWD signals out to several tens of Mpc. Using SN Ia observations to estimate the corresponding DWD merger rate (assuming that all SN Ia are produced by DWDs), we found in section 3.2.2 that about 20 DWDs producing SN Ia would be observed in 10 years with LGWA. This would be enough detections to make a meaningful comparison of rates and infer a possible DD progenitor scenario.
Interestingly enough, the simultaneous detection of a “peculiar” SN Ia in both GW and EM domains is expected for thermonuclear events arising from WD-WD systems originating from a CO-WD and a He-WD. These objects are anticipated to explode via the detonation of a thin He-layer piled-up on the CO-WD through a stable mass transfer that is driven by GW emission with maximum expected frequency of several mHz (see [847] for details). Multi-messenger astronomy could help determining the contribution of double-detonating WDs to the observed SNe Ia rate and, more broadly, allow one to infer the chirp masses for some of the transient, hence constraining the SN progenitor.
@@
Following the gravitational collapse of a massive star ($M\geq 8M_{\odot}$), the inner core rebounds and collides with the infalling stellar material producing a shock wave (see [848] for a general review). The shock wave tends to stall within the infalling material. Extensive theoretical work done in recent years has demonstrated that it is the neutrino heating coming from the accreting proto-NS (coupled to the neutrino-driven turbulent convection between the shock cavity and the stellar surface) that re-powers the shock and leads to subsequent explosion ([849, 338, 329] and references therein). The shock break-out, i.e., the moment when the blast wave reaches the stellar surface, occurs minutes to days post collapse, and shortly after starts emitting electromagnetic radiation, appearing as a SN. The star is destroyed leaving behind a compact remnant.
@@ 3.3.3 Tidal Disruption Events
Indeed, ZTF EM follow-up searches of IceCube alerts identified TDE AT2019dsg [907] as a possible optical counterpart of a very high energy neutrino ($\sim$0.2 PeV, IceCube event IC191001A), which represents the first multi-messenger observation of a TDE. Later, other two TDE candidates have been associated with two high-energy neutrino events: AT2019fdr (IceCube event IC200530A [908]) and AT2019aalc (IceCube event IC191119A [909]). Interestingly, all the three neutrino-emitting TDEs discovered so far show some common observational features, such as the delayed neutrino signal ($\sim$100 days after the BB peak), a delayed strong dust echoes in the IR and the X-ray detection. This can be ascribed to the physical properties of the system after the star disruption and raise important question on the mechanisms behind the neutrino production itself [910]. However, it should be noted that recent reanalyses of IceCube data suggest that the coincidences between TDE and neutrino observations are consistent with background [911].
@@ 3.3.4 Intermediate and Massive Black Hole Binaries
Where to look for electromagnetic candidates for LGWA emitters ? – The tight correlations observed between the host-galaxy properties and the BH mass indicate that associated EM signals produced from the more massive BHs detected by LGWA, i.e., upwards of 105 – 106 M, should be mainly searched for in dwarf galaxies. Even more challenging is the case of lower-mass BHs (IMBHs). Indeed, besides some indications from GWs [945, 366], any conclusive electromagnetic evidence for the existence of even a single IMBH is still missing. Among the best candidates for accreting IMBHs, we can include ultra-luminous X-ray sources (ULXs; e.g., [946, 947]). ULXs are defined as point-like, off-nuclear (i.e., sources showing offsets from the dynamical center of their hosts) X-ray sources in nearby galaxies; they are characterized by X-ray luminosity above the Eddington limit for a stellar BH (LX$>$1039 erg s-1). In particular, the still-limited (in number) high-mass end of ULX distribution could be powered by IMBHs. ULXs are largely associated with star-forming regions in spiral and irregular galaxies and possibly blue compact dwarf galaxies; however, cases where they trace the old stellar population in elliptical galaxies are also present (see [948]). Since low-mass BHs in shallow potentials (as, e.g., those of dwarf galaxies) are likely to never settle at the center of their galaxies [949], we may expect that the location of IMBHs is consistent with the offset position of ULXs. Another astrophysical environment in which low-mass BHs may form is that of globular clusters [920]. As dense stellar clusters evolve, the compact remnants sink to the center of the cluster and segregate in mass. Although old globular clusters are not known for their large gas content, dynamical interactions in the densest regions may convey material (possibly originated by gas lost by stars during their evolutionary phases). If the gas available in such systems is sufficient to fuel efficient accretion, electromagnetic signatures are likely to be expected. Wandering IMBHs in the halo of galaxies are also expected as a result of BH-BH encounters in globular clusters that can eject IMBHs. These objects can also represent one of the consequences of the galaxy merging process if seed BHs were common in the first galaxies that merged to form today’s massive galaxies.
@@ 3.3.5 Extreme/Intermediate Mass-ratio Inspirals
So far the signal observed from QPEs is restricted to the soft X-ray band ($E<2$ keV) and the abundance rates of these transients are low ($\approx 0.6\times 10^{-6}\,$pc-3; [1033]). Therefore, for the purpose of QPE discovery, a soft X-ray telescope with a large effective area and a large field of view is desirable. The WFI onboard Athena may discover $\approx 2$ QPE source per 15-square-degree survey [1033]. WFI will also be able to study spectral and timing properties of QPEs deeper in redshift, up to $z\sim 0.5$ for the most common QPE sources peaking at $\sim 10^{42}$ erg s-1 and up to $z\sim 1.3$ for the much rarer ones peaking at $\sim 10^{43}$ erg s-1. The large effective area and high spectral resolution of the X-IFU onboard Athena will also allow to assess the presence of ionized outflowing/inflowing gas before, during, and after QPEs.